OpenStudy (kidtalk96):

please solve. 3x + 5y = 1 4x +2y = -8 i cant figure out how to get my answer to these kinds of problems. please help.

OpenStudy (anonymous):

Combine both equations (by adding or subtracting them) and at the same time, eliminate one of the variables. Then you have one equation with one variable, solve it and you have one of the solutions. Next use that on either equation and you get the second solution.

OpenStudy (anonymous):

You do that by having the same value for one variable on both equations. So lets say you multiply the first equation by 4 and the second equation by 3 what do you get?

OpenStudy (anonymous):

4( 3x+5y=1) is: 12x+50y=4 And: 3(4x+2y=-8) is: 12x+6y=-24

OpenStudy (anonymous):

Now you have the same coefficient for x on both equations, Now take one equation and subtract it from the other. 12x+20y=4 minus CORRECTION! IT IS 20y NOT 50y !!!! 12x+6y=-24 You get the x eliminate from the equation and now you have only y as the variable: 14y=28 or y=28/14 y=2

OpenStudy (anonymous):

Now, just substitute y=2 on either equation to get the value for x: 3x+5y=1 substituting y=2: 3x+5(2)=1 or 3x+10=1 next 3x=-9 so x=-3

OpenStudy (anonymous):

3x+5y=1}x2 6x+10y=2 (1) -14x=42=>x=-3 => => (1)-(2) 3(-3)+5y=1=>y=2 4x+2y=-8}x5 20x+10y=-40 (2)